OpenStudy (anonymous):

11 and 7/8 - 4 and 5/12

OpenStudy (anonymous):

do you mean \[11+\frac{7}{8}-4+\frac{5}{12}\]

OpenStudy (anonymous):

Its 11 7/8 11 is the whole number of that fraction and 4 is the whole number of 5/12

OpenStudy (anonymous):

\[11+\frac{7}{8} - 4 - \frac{5}{12}\] \[11-4 = 7\] \[7 + \frac{7}{8}-\frac{5}{12}\] common denominator of 24. simplify

OpenStudy (anonymous):

7 1/2?